OpenStudy (anonymous):

Ron borrows $ 8000 at 12% per annum simple interest and Mike borrows $9100 at 10% per annum simple interest. In how many years will their borrowed amounts(debt) be equal ? a)18 b)20 c)22 c)24

OpenStudy (anonymous):

we can use Arithmetic Sequence and its Sum...

OpenStudy (anonymous):

how ?

OpenStudy (anonymous):

the 1st item for Ron, will be the 1st year of his debt, that is... p1=8000(1.12)=8960 p2=8960+960=9920 p3=8960+960(2)=10880 our common d=960=8000(0.12)=the interest per annum...

OpenStudy (anonymous):

so the sum of debt for Roy (let it be SRoy) at t years would be... \[SRoy=\frac{t}{2}(2(8960)+960(t-1))\]

OpenStudy (anonymous):

we'll do the same for Mike... and equate the sum of their debts to solve for t years...

OpenStudy (anonymous):

got the idea @moli1993 ?

OpenStudy (anonymous):

the sequence for Mike will be as follows: p0=9100 p1=9100(1.10)=10010 p2=9100(1.20)=10010+910=10920 the common d=910=9100(0.10)=interest for Mike...

OpenStudy (anonymous):

same formula applies for the sum of Mike's debt for t years...\[SMike=\frac{t}{2}(2(10010)+910(t-1))\]

OpenStudy (anonymous):

Equating SRoy=SMike will allow us to solve t years...

OpenStudy (anonymous):

ohk u r applying arithmetic progression i.e. sum=n/2(2a+(n-1)d)

OpenStudy (anonymous):

yup...

OpenStudy (anonymous):

that's the way i see it...

OpenStudy (anonymous):

ohh cool but i cudnt understand why are u taking 8960 instead of 8000 in the formula ? and vice versa for 9100?

OpenStudy (anonymous):

that's the start of their payment if they'll pay for 1st year of their borrowed money...

OpenStudy (anonymous):

ummm i calculated but the answer is not matching with any of the options :(

OpenStudy (anonymous):

it's coming 43

OpenStudy (anonymous):

maybe the others have different idea...

OpenStudy (anonymous):

i'm thinking also it it's a geometric progression but can't see a common ratio....

OpenStudy (anonymous):

can u do it using the formula simple interest= principle*rate*time/100

OpenStudy (anonymous):

time will just cancel once you equate the two...

OpenStudy (anonymous):

@ganeshie8 can u help us out !

OpenStudy (anonymous):

yes time has cancelled out, the inner one is remaining and t=43 is coming on solving this equation

OpenStudy (anonymous):

hmmm... let us try equation of a line... simple interest is a linear function... let us use the slope-intercept formula...\[y=mx+b\]Let's have some definitions:\[y=S(t)=principal+interest\]\[m=principal\times rate\]\[b=principal\]\[t=time~(years)\]The equation for Roy's scheme will be: \[S(t)=960t+8000~~~~(1)\]For Mike's scheme it will be:\[S(t)=910t+9100~~~~(2)\]Equating (1) and (2) will yield\[960t+8000=910t+9100\]\[(960-910)t=9100-8000\]\[50t=1100\]\[t=\frac{1100}{50}\]\[t=22~years\]

OpenStudy (anonymous):

... so it is letter c... \(\ddot\smile\)